# Simple Policy Optimization

### v7

@@ 5 Theoretical Results
In this section, we provide some theoretical insights of the differences between PPO and SPO, demonstrate that SPO can be more effective in constraining probability ratios.
@@ 6.1 Comparing Algorithms
Our implementation of SPO is compared against PPO-Clip (Schulman et al., 2017), PPO-Penalty (Schulman et al., 2017), SPU (Vuong et al., 2018), PPO-RB (Wang et al., 2020), TR-PPO (Wang et al., 2020), TR-PPO-RB (Wang et al., 2020), and RPO (Gan et al., 2024) in MuJoCo benchmark. We compute the algorithm’s performance across ten separate runs with different random seeds. In addition, we emphasize that in all comparative experiments involving the same settings for SPO and PPO, the only modification in SPO is replacing the PPO’s objective with (16), no further code-level tuning is applied to SPO, highlighting its simplicity and effectiveness.
@@
Table 1: Average return of PPO and SPO in the last 10% training steps across five separate runs with different random seeds, with their maximum ratio deviation of the entire training process.
@@ 6.2 Scaling Policy Network
Figure 6: Training curves of SPO using ResNet-18 as the encoder compared to the PPO and SPO with original CNN. The mean and standard deviation are shown across 3 random seeds.
@@
We also trained the ResNet-18111Since Bjorck et al. (2021) demonstrated that batch normalization is harmful to RL, we removed batch normalization and adjusted ResNet-18 for input and output from our implementation. (He et al., 2016) as the encoder on the Atari 2600 benchmark, the results are shown in Figure 6. As the network’s capacity increases, the performance of SPO is significantly improved. Moreover, SPO can still maintain a good probability ratio constraint, thereby benefiting from the theoretical lower bound (10). In contrast, it is challenging to train large neural networks with PPO because the probability ratio can not be controlled during training, even employing a smaller $\epsilon=0.1$.
@@ Appendix C More Results
Figure 9: Aggregate metrics on MuJoCo-v4 with 95% CIs based on 6 environments, comparing PPO and SPO with different policy network layers and mini-batch size using PPO-normalized score.

### v8

@@ 5 Theoretical Results
In this section, we provide some theoretical insights of the differences between PPO and SPO, demonstrating that SPO can be more effective in constraining probability ratios.
@@ 6.1 Comparing Algorithms
Our implementation of SPO is compared against PPO-Clip (Schulman et al., 2017), PPO-Penalty (Schulman et al., 2017), SPU (Vuong et al., 2018), PPO-RB (Wang et al., 2020), TR-PPO (Wang et al., 2020), TR-PPO-RB (Wang et al., 2020), and RPO (Gan et al., 2024) in MuJoCo benchmark. We compute the algorithm’s performance across ten separate runs with different random seeds. In addition, we emphasize that in all comparative experiments involving the same settings for SPO and PPO, the only modification in SPO is replacing the PPO’s objective with (16), no further code-level tuning is applied to SPO, highlighting its simplicity and efficiency.
@@
Table 1: Average return of PPO and SPO in the last 10% training steps across 5 separate runs with different random seeds, with their maximum ratio deviation during the entire training process.
@@ 6.2 Scaling Policy Network
Figure 6: Training curves of SPO using ResNet-18 as the encoder compared to the PPO and SPO using original CNN. The mean and standard deviation are shown across 3 random seeds, and the red dashed line shows $\epsilon=0.2$.
@@
We also trained the ResNet-18111Since Bhatt et al. (2019) demonstrated that batch normalization is harmful to RL, we removed batch normalization and adjusted ResNet-18 for input and output from our implementation. (He et al., 2016) as the encoder on the Atari 2600 benchmark, the results are shown in Figure 6. As the network’s capacity increases, the performance of SPO is significantly improved. Moreover, SPO can still maintain a good probability ratio constraint, thereby benefiting from the theoretical lower bound (10). In contrast, it is challenging to train large neural networks with PPO because the probability ratio cannot be controlled during training, even employing a smaller $\epsilon=0.1$.
@@ 7 Conclusion
Acknowledgements
We thank the anonymous ICLR and ICML reviewers for their insightful and constructive comments, as well as our collaborators for their valuable suggestions.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix C More Results
Figure 9: Aggregate metrics on MuJoCo-v4 with 95% CIs based on 6 environments, comparing PPO and SPO with different policy network layers and mini-batch sizes using PPO-normalized score.